OpenStudy (jiteshmeghwal9):

hey guys if anybody knows to find the no. of polynomials then it's ok & if anyone don't knows then have a look:)

OpenStudy (jiteshmeghwal9):

No. of diagonals in a polygon if it has 'n' vertices:- \[no. \space of \space diagonals=\frac{n(n-3)}{2}\]

OpenStudy (jiteshmeghwal9):

Example: Polygon with 22 vertices. No. of diagonal=\[\frac{n(n-3)}{2}\]\[22(22-3)\over2\]\[22 \times 19\over2\]\[11 \times 19=209\]

OpenStudy (shubhamsrg):

and derivation of that?

Parth (parthkohli):

Haha! I just went through this formula in my class :)

Parth (parthkohli):

Also, please correct your question to "no. of diagonals." You've written - 'no. of polynomials'. :P

OpenStudy (jiteshmeghwal9):

sorry ! @ParthKohli

Parth (parthkohli):

Nah. This was very helpful for curious people like me =)

OpenStudy (jiteshmeghwal9):

thanx @ParthKohli

OpenStudy (jiteshmeghwal9):

the derivation is done through example @shubhamsrg

OpenStudy (shubhamsrg):

lol.. i meant how do you get that formulla for a n sided figure ?

OpenStudy (shubhamsrg):

the derivation, the proof!! ??

OpenStudy (jiteshmeghwal9):

i study about this formula through a site there they don't taught me the proof

OpenStudy (jiteshmeghwal9):

didn't*

OpenStudy (jiteshmeghwal9):

@shubhamsrg

OpenStudy (shubhamsrg):

hmm..

OpenStudy (foolaroundmath):

It can be derived if you know a bit of Combinatorics.  Number of ways to select any two corners = \(^{n}C_{2}\) Now among all selections of any two corners, it wont form a diagonal if the two points are adjacent to each other. No. of ways of selecting two adjacent points = \(n\) (Why?) So number of diagonals = \(^{n}C_{2} - n\)\[= \frac{n(n-1)}{2}-n= \frac{n(n-3)}{2}\]